OpenStudy (anonymous):

Find the value of x ..... 2x=7+4

OpenStudy (anonymous):

add 7+4 2x=11 divide 2: x=11/2

OpenStudy (anonymous):

u can get a mix number n a fraction too or u leave it like dat

OpenStudy (anonymous):

@jamesumner93

OpenStudy (anonymous):

add 7 and 4 then divide by 2 so x=11/2 or 5 1/2 :)

OpenStudy (anonymous):

okay thanks @Hally

OpenStudy (anonymous):

U're welcome! :)